OpenStudy (anonymous):

An arrow hits the target at 17.0m/s and penetrates for 0.0085s at an acceleration of -1550m/s^2. The target's magnitude has an acceleration of 450 m/s^2 How far does the arrow penetrate?

OpenStudy (anonymous):

we use v^2=u^2+2as where v=final velocity =0; initial velocity =17m/s and a=-1550m/s^2 then u^2=2as and s= distance=u^2/2a s=.093m